OpenStudy (anonymous):

How many grams of Al2(Cr)4)3 can be produced from 125 mL of a 0.150 Molarity K2CrO4 solution? Balanced equation: 2 Al(NO3)3(aq) + 3K2CrO4(aq) = Al2(CrO4)3(s) + 6KNO3(aq) Molar Mass: 213 194 402 111

sam (.sam.):

Ok so do you know where to start?

sam (.sam.):

We first have to find the number of moles of \(K_2CrO_4\) solution, then use that to switch using mol to mol conversion.

OpenStudy (anonymous):

ok so .125L * 0.150 M = .01875 moles K2CrO4

sam (.sam.):

Now you can convert that using mol to mol conversion using the coefficients on the reaction equation, \[0.01875mol~K_2CrO_4 \times \frac{1~mol~Al_2(CrO_4)_3}{3~mol~K_2CrO_4 }=\]

sam (.sam.):

After that, you'll have \(~moles~of~Al_2(CrO_4)_3\). Then multiply it by the relative molecular mass to get the mass of \(Al_2(CrO_4)_3\)

OpenStudy (anonymous):

okay awesome!! i got it now! thanks.. it can get confusing when i can't find an example in my notes close enough to figure out all the steps to go by

sam (.sam.):

Oh you're glad you've got one here :)